OpenStudy (anonymous):

what's the domain and range of y = -2/x? is the domain and range of a reciprocal function always (- infinity, 0) (-infinity,infinity)?

OpenStudy (e.mccormick):

The domain can not make the fraction invalid. so if it is just over x, \(x^2\) or something lke thatm then the domain will avoid 0, but whjat about \(\frac{1}{x+2}\)? The problem 0 there is at x=2.

OpenStudy (e.mccormick):

Oops. Lissed a - in my excample.

OpenStudy (e.mccormick):

And the range is what are the valid outputs of the function. In this casem how would it ever be equal to 0? So the range also avoids 0 in this situation.

OpenStudy (e.mccormick):

Here is a graph to help you visualize this. Remember, valid x values are the domain and valid y values are the range. What you put above, Domain \(= (- \infty, 0)\) Range \(= (-\infty,\infty)\), is not correct.